### v1

# A Bayesian approach for fitting semi-Markov mixture models of cancer latency to individual-level data

Abstract
Joint senior authors
Multi-state models of cancer natural history are widely used for designing and evaluating cancer early detection strategies. Calibrating such models against longitudinal data from screened cohorts is challenging, especially when fitting non-Markovian mixture models against individual-level data. Here, we consider a family of semi-Markov mixture models of cancer natural history introduce an efficient data-augmented Markov chain Monte Carlo sampling algorithm for fitting these models to individual-level screening and cancer diagnosis histories. Our fully Bayesian approach supports rigorous uncertainty quantification and model selection through leave-one-out cross-validation, and it enables the estimation of screening-related overdiagnosis rates. We demonstrate the effectiveness of our approach using synthetic data, showing that the sampling algorithm efficiently explores the joint posterior distribution of model parameters and latent variables. Finally, we apply our method to data from the US Breast Cancer Surveillance Consortium and estimate the extent of breast cancer overdiagnosis associated with mammography screening. The sampler and model comparison method are available in the R package baclava.
Keywords Cancer natural history  $\cdot$ Cancer indolence  $\cdot$ Overdiagnosis  $\cdot$ Data augmentation  $\cdot$ Markov chain Monte Carlo
1 Introduction
Cancer screening aims to detect early signs or markers of cancer before symptoms arise, facilitating timely intervention for improved patient outcomes. The efficacy of screening depends on the duration of the pre-clinical phase, during which the tumor is asymptomatic but screen-detectable. If the latency period of a tumor is too short, even frequent screening does not significantly advance the time of diagnosis. Conversely, a long latency period can result in overdiagnosis—the diagnosis and unnecessary treatment of slowly growing tumors that would not have caused symptoms or other harm in the person’s remaining lifetime (Welch and Black, 2010; Duffy and Parmar, 2013). Understanding the duration of tumor latency is thus crucial for assessing the balance between the benefits and harms of a screening test.
Estimating tumor latency from cancer incidence data is challenging, because the onset of pre-clinical disease in healthy individuals is not directly observable. In unscreened populations, only the time of clinical diagnosis, which coincides with the end of the tumor latency period, is known. In this case, we cannot distinguish the time intervals before and after pre-clinical onset. In other words, the period of tumor latency is not identifiable. Fortunately, screening methods that utilize biological or imaging markers offer a selective insight into the latent disease state, thus enabling the identification of tumor latency.
There is a rich history on estimating cancer latency from screening data using multi-state models that dates back to the 1980s. These models range from simple three-state models (Day and Walter, 1984; Shen and Zelen, 1999)—in which individuals are healthy (disease-free), pre-clinical (with an asymptomatic screen-detectable cancer) or clinical (with a symptomatic cancer)—to more complex models that seek to emulate the biological complexity of cancer progression (Rutter and Savarino, 2010; Seigneurin et al., 2011). Typically, the underlying stochastic processes of cancer natural history models are assumed to be Markovian for mathematical tractability. Under this assumption, the unobserved onset times can be integrated out analytically, giving a closed-form expression for the likelihood of observed data. This enables maximum likelihood estimation (Olsen et al., 2006; Wu et al., 2018) and Bayesian inference (Duffy et al., 2005; Ryser et al., 2022). However, relaxing the Markov assumption of state transitions often leads to a departure from the simpler mathematical structures, and the calibration of more complex non-Markovian models is usually performed using simulation-based or “likelihood-free” approximation methods, such as approximate Bayesian computation (Rutter et al., 2009; Alagoz et al., 2018; Bondi et al., 2023) or pseudo-likelihood methods (Moore et al., 2001).
In this paper, we present a method for exact Bayesian inference of cancer natural history using a flexible family of semi-Markov mixture models that account for both progressive and non-progressive (indolent) disease states. The proposed approach accommodates individual-level screening and cancer diagnosis histories, and, because it does not rely on model-based simulation, is scalable to large cohorts of screened individuals. By relaxing the Markov assumption, we may then flexibly specify the hazard function, which in turn leads to improved model realism (Hsieh et al., 2002; Cheung et al., 2022). Since the semi-Markov models do not admit a closed form expression for the observed data likelihood, we conduct inference using data augmentation (Tanner and Wong, 1987). This strategy augments the observed data with latent variables in such a way that the resulting complete-data likelihood is amenable to iterative sampling. Executing this idea in the context of cancer natural history modeling is nontrivial because there are twice as many latent variables as the number of individuals in the study, and cancer screening cohorts often comprise tens or hundreds of thousands of individuals. The resulting joint posterior distribution is complex and high-dimensional, rendering standard rejection or importance sampling approaches impractical. Moreover, gradient-based approaches, such as Hamiltonian Monte Carlo, which are well-suited for high-dimensional problems (Girolami and Calderhead, 2011), are not applicable to our family of models due to the presence of discrete latent variables, and of discontinuities in the likelihood from the discrete screening outcomes.
Instead, our data-augmented Markov chain Monte Carlo (DA-MCMC) sampler efficiently explores the joint posterior distribution of the model parameters together with the latent variables, enjoying sound theoretical guarantees. This fully Bayesian approach automatically accounts for the uncertainty in the exact onset time of pre-clinical cancers, and allows the estimation of any relevant function of the parameters such as the overdiagnosis rate through their posterior distribution. In addition, because the complete-data likelihood has a simple closed form, we can rigorously compare and select models using an extension of approximate leave-one-out cross-validation (Gelfand et al., 1992; Gelfand, 1995; Vehtari et al., 2017) to latent data models.
The paper is structured as follows. In Section 2, we introduce the underlying family of semi-Markov mixture models of cancer natural history; in Section 3, we describe the inferential framework; and, in Section 4, we detail the construction of the DA-MCMC algorithm. We examine the mixing properties of the sampler through simulation studies in Section 5 and apply it to estimate the extent of overdiagnosis in a real-world cohort of women undergoing mammography screening in Section 6. The DA-MCMC sampling algorithm is publicly available in the R package baclava, and R code for reproducing the results present in this paper is available on GitHub (https://github.com/rmorsomme/baclava-manuscript).
2 Semi-Markov mixture model of cancer progression
Healthy ($H$)Pre-clinical cancer ($P$) Clinical cancer ($C$)progressiveindolentY1.0e0]ptY1.0e0]pt
(a) Compartmental model for cancer progression with a mixture of progressive and indolent pre-clinical cancers.
[…]
(b) Illustration of possible individual clinical trajectories through the compartmental model.
Figure 1: Panel (a) depicts the compartmental model for the progression of cancer. Panel (b) illustrates possible individual trajectories through the compartmental model. Individuals start in compartment $H$. Only transitions from $H$ to $P$ and from $P$ to $C$ are possible, with compartment $P$ being an absorbing state for indolent individuals and compartment $C$ being absorbing for all others. In Panel (b), individual 1 does not develop pre-clinical cancer before the end of the observation period and, therefore, remains in $H$ until that time. Individual 2 develops pre-clinical cancer at age $\tau^{HP}_{2}$, at which time they transition to $P$, but they do not develop clinical cancer before the end of the observation period and, therefore, remain in $P$. Note that the pre-clinical cancer of individual 2 may be either indolent, in which case $\tau^{PC}_{2}=\infty$, or progressive with a clinical onset time after the end of the observation period. Individual 3 develops pre-clinical cancer at age $\tau^{HP}_{3}$ and transitions to state $P$. At age $\tau^{PC}_{3}$, they develop clinical cancer and thus transition to $C$. As $C$ is absorbing, the observation period ends with their transition time into $C$.
We model the dynamics of cancer progression as a multi-state mixture model. The model, depicted in Figure 1(a), consists of the three compartments healthy ($H$), pre-clinical cancer ($P$) and clinical cancer ($C$). Following Ryser et al. (2019), pre-clinical cancers are modeled as a mixture of progressive and indolent tumors. All individuals start in the healthy state $H$, and transition to a state of pre-clinical screen-detectable cancer $P$ after a randomly distributed time. Once in $P$, individuals with an indolent cancer stay in this compartment indefinitely, while individuals with a progressive cancer eventually transition to $C$ when their cancer becomes symptomatic.
Several early works consider purely progressive models, in which all pre-clinical tumors progress to clinical disease after a finite time (Day and Walter, 1984; Shen and Zelen, 1999, 2001). Allowing a proportion of pre-clinical cancers to be non-progressive is particularly important for slowly growing or indolent disease entities, such as low-risk prostate cancers or in situ breast cancers (Baker et al., 2014). Despite the challenges of fitting such mixture models in practice (Ryser et al., 2019), these models have gained in popularity, especially in the study of breast cancer progression (Chen et al., 1996; Duffy et al., 2005; Olsen et al., 2006; Seigneurin et al., 2011; Alagoz et al., 2018; Wu et al., 2018; Ryser et al., 2022).
To formalize the model structure, we let the starting age $t_{0}\geq 0$ correspond to the age at which individuals become susceptible to developing pre-clinical cancer, so that the hazard rate of pre-clinical cancer is assumed to be zero before $t_{0}$. Next, let $\tau^{HP}$ be the age at pre-clinical onset (the transition from $H$ to $P$) and let $\tau^{PC}$ be the age at clinical onset (the transition from $P$ to $C$); by definition, $t_{0}\leq\tau^{HP}\leq\tau^{PC}$. When a tumor is indolent, we set $\tau^{PC}=\infty$, reflecting the fact that indolent cancers never progress to the clinical stage. Figure 1(b) illustrates the evolution of three individuals throughout this multi-state model.
The associated stochastic process is characterized by the probabilistic mechanisms governing the transitions between the compartments. Denote the waiting time until pre-clinical cancer with $\sigma^{H}=\tau^{HP}-t_{0}$ and the pre-clinical sojourn time with $\sigma^{P}=\tau^{PC}-\tau^{HP}$, corresponding to the interval between the onset of pre-clinical cancer and the diagnosis with symptomatic disease. An individual with an indolent tumor has $\sigma^{P}=\infty$. We assume that
$\sigma^{H}\sim\textsf{F}_{\textsf{H}},\quad\sigma^{P}\sim\textsf{F}_{\textsf{P}},$ (1)
independently, where $\textsf{F}_{\textsf{H}}$ is a distribution on the positive line, and $\textsf{F}_{\textsf{P}}$ corresponds to the mixture cure model (Berkson and Gage, 1952; Peng and Taylor, 2014),
$\textsf{F}_{\textsf{P}}=\psi\delta_{\infty}+(1-\psi)\textsf{F}_{\textsf{prog}}.$ (2)
with a point mass at $\infty$ for indolent cancers and a density on the positive line for progressive cancers. Here, $\psi\in[0,1]$ is the proportion of indolent cancers and $\delta_{\infty}$ the Dirac distribution with point mass at $\infty$. Distribution (2) implies that a proportion $\psi$ of tumors are indolent with infinite sojourn times while the sojourn time of progressive tumors follows distribution $\textsf{F}_{\textsf{prog}}$.
Our inferential framework will apply to arbitrary choices of parametric distributions on the positive line for $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$, resulting in a flexible semi-Markovian formulation. We write $\theta$ to refer to the collective parameters of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$. This formulation encompasses existing models as special cases: for example, letting $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ be exponential distributions results in the Markov mixture model of Seigneurin et al. (2011), and $\psi=0$ recovers the purely progressive model of Shen and Zelen (1999, 2001). Introducing the indolence indicator $I$, which equals $1$ if the cancer is indolent and 00 otherwise, we can also express the distribution (2) hierarchically as
$I$ $\sim\textsf{Ber}(\psi),$ (3)
$\sigma^{P}\mid I$$\sim\begin{cases}\delta_{\infty},&\quad I=1\\ \textsf{F}_{\textsf{prog}},&\quad I=0\end{cases},$ (4)
where $\textsf{Ber}(p)$ denotes the Bernoulli distribution with probability parameter $p$.
3 Likelihood and inferential methodology
3.1 Likelihood of complete data
Consider a sample of $n$ individuals that evolve independently, and let the subscript $i=1,\dots,n$ refer to quantities related to the $i$th individual. We refer to the data for individual $i$ as $X_{i}=\left(I_{i},\tau^{HP}_{i},\tau^{PC}_{i}\right)$ and write $\mathbf{X}=(X_{1},\dots,X_{n})$ for the data of the entire population. By independence between the individuals, the likelihood of the process described by equations (1)-(4) factorizes as
$\tilde{L}(\theta,\psi;\mathbf{X})=\prod_{i=1}^{n}\tilde{L}_{i}(\theta,\psi;X_{i})$ (5)
where
$\tilde{L}_{i}(\theta,\psi;X_{i})=\begin{cases}S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right),&\quad c_{i}<\tau^{HP}_{i}\\ f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)f_{\textsf{I}}(I_{i};\psi)S_{\textsf{prog}}\left(c_{i}-\tau^{HP}_{i};\theta\right)^{1-I_{i}},&\quad\tau^{HP}_{i}\leq c_{i}<\tau^{PC}_{i}\\ f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)(1-\psi)f_{\textsf{prog}}\left(\sigma^{P}_{i};\theta\right),&\quad\tau^{PC}_{i}\leq c_{i}\end{cases}.$ (6)
is the likelihood of person $i$ observed until their censoring age $c_{i}$. Here, $f_{\cdot}$ and $S_{\cdot}$ are the density and survivial function of the corresponding distribution $\textsf{F}_{\cdot}$ and $f_{\textsf{I}}$ is the probability mass function of the Bernoulli distribution (3).
We can understand the individual likelihood (6) as follows. The contribution of a person $i$ to the likelihood $\tilde{L}(\theta;\mathbf{X})$ depends on which compartment they are in at age $c_{i}$. If individual $i$ does not develop pre-clinical cancer by the censoring age $\left(c_{i}<\tau^{HP}_{i}\right)$, then their contribution is the survival function $S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right)$ of $\textsf{F}_{\textsf{H}}$. If they develop pre-clinical cancer but not clinical cancer by the censoring time $\left(\tau^{HP}_{i}\leq c_{i}<\tau^{PC}_{i}\right)$, then their contribution consists of the product of the density $f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)$ of $\textsf{F}_{\textsf{H}}$, the Bernoulli probability mass function of the indolence indicator $f_{\textsf{I}}(I_{i};\psi)$, and $S_{\textsf{prog}}\left(c_{i}-\tau^{HP}_{i};\theta\right)^{1-I_{i}}$ which is either the constant function $1$ for indolent cases $(I_{i}=1)$ or the survival function of $\textsf{F}_{\textsf{prog}}$ for progressive cases ($I_{i}=0$). Finally, if the individual develops clinical cancer by the censoring time $\left(\tau^{PC}_{i}\leq c_{i}\right)$, then their contribution is the product of $f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)$, the probability $1-\psi$ of a progressive cancer, and the density $f_{\textsf{prog}}\left(\sigma^{P}_{i};\theta\right)$ of distribution $\textsf{F}_{\textsf{prog}}$.
3.2 Partially observed data
Age
(in years)
[…]
(a) Censored individual.
Age
(in years)
[…]
(b) Screen-detected individual.
Age
(in years)
[…]
(c) Interval-detected individual
Individual Screen age Screen outcome
$1$ $40$ 00
$1$ $45$ 00
$2$ $40$ 00
$2$ $45$ $1$
$3$ $40$ 00
$3$ $45$ 00
(d) Observed screen data for the 3 individuals corresponding to Figures 2(a)-2(c).
Individual $t^{PC}_{i}$ Censoring age
$1$ $46.5$ $46.5$
$2$ $45$ $45$
$3$ $46$
(e) Observed right-censored transition age into $C$, $t^{PC}_{i}$, and censoring age for the 3 individuals corresponding to Figures 2(a)-2(c).
Figure 2: Observed data for the individuals depicted in Figure 1(b). Panels (a-c): diagram displaying the trajectories and screen results. Panel (d): screen results. Panel (e): right-censored onset age of clinical cancer. In Panel (a), the individual does not develop pre-clinical cancer before their censoring age. They are in $H$ at the two screen times, the two screens are therefore negative. They are censored at age $46.5$, $1.5$ years after their last screen. In Panel (b), the individual develops a pre-clinical before age $40$. They are in $P$ when the two screens take place. The first screen is a false negative, and the second is positive. They are censored at age $45$, their age at the positive screen. In Panel (c), the individual develops a pre-clinical cancer between the first and second screen and a clinical cancer at $46$. Their screen at age $45$ is a false negative. The onset time of their clinical cancer at age $46$ is observed.
We now turn our attention to real-world data sources, such as prospective screening trials and observational screening cohorts, In such settings, the data consist of a series of screens and associated results, as well as clinical cancer diagnoses. Let $n_{i}$ denote the number of screens available for person $i$, with their age and outcome at each of these screening times contained in the vectors $\left(t^{i}_{1},\dots t^{i}_{n_{i}}\right)$ and $\left(O^{i}_{1},\dots O^{i}_{n_{i}}\right)\in\{0,1\}^{n_{i}}$, respectively, where $t_{1}^{i}=e_{i}$ corresponds to the age at study entry. These data are complemented by the time $t^{PC}_{i}=\min\left\{c_{i},\tau^{PC}_{i}\right\}$, which corresponds to either the age of diagnosis with a screen-detected cancer, the age of diagnosis with a clinical cancer, or to the age of censoring otherwise.
We assume that the screening ages are non-informative, that is, that they are ignorable or independent of the disease process and of the model parameters. Under this assumption, we may condition on them in our analysis (Gelman et al., 2013). We therefore denote the observed variables of individual $i$ with $Y_{i}=\left(O^{i}_{1},\dots,O^{i}_{n_{i}},t^{PC}_{i}\right)$ and condition on the ignorable screen ages $(t^{i}_{1},\dots,t^{i}_{n_{i}})$. We further write $\mathbf{Y}=(Y_{1},\dots,Y_{n})$. Tables 2(d) and 2(e) present the observed data of a sample of three individuals for illustration.
The screening outcomes are modeled as mutually independent Bernoulli random variables
$O^{i}_{j}\sim\textsf{Ber}\left(p^{i}_{j}\right),\quad j=1,\dots,n_{i},\quad i=1,\dots,n,\qquad\text{where}\quad p^{i}_{j}=\begin{cases}0,\quad t^{i}_{j}<\tau^{HP}_{i}\\ \beta,\quad\tau^{HP}_{i}\leq t^{i}_{j}<\tau^{PC}_{i}\\ \end{cases}.$ (7)
Equation (7) indicates that the probability of positive screen $p^{i}_{j}$ is 00 when a person is in the $H$ compartment, while it becomes $\beta$ when they are in $P$; we see the parameter $\beta\in[0,1]$ is interpreted as the screening sensitivity and is not assumed to be known. Upon a cancer diagnosis—whether clinical or screen-detected—an individual receives treatment and leaves the screening program, implying that any positive screen is the last screen of the individual. Consequently, $p^{i}_{j}$ need not be defined when $t^{i}_{j}>\tau^{PC}_{i}$.
The observed data partition the individuals into three groups. The censored group consists of individuals with no positive screens and who do not develop a clinical cancer before their censoring age. These individuals either remain in $H$ until their censoring age, or transition into $P$ and remain there until their censoring age; any screens conducted when they are in $P$ are false negatives (see Figure 2(a)). The screen-detected group includes individuals who are in state $P$ at the time of the positive screen (see Figure 2(b)). Finally, the interval-detected group contains those with no positive screen but whose cancer reaches the clinical stage before their censoring age (see Figure 2(c)). These individuals may or may not have false negative screens.
3.3 Bayesian inference via data-augmented Markov chain Monte Carlo
We now describe how to conduct Bayesian inference to fit the semi-Markov model defined by Equations (1)-(4) to censored data $\mathbf{Y}$. By Bayes Theorem, $\pi(\theta,\psi,\beta\mid\mathbf{Y})\propto\pi(\theta,\psi,\beta)L(\theta,\psi,\beta;\mathbf{Y})$ where $\pi(\theta,\psi,\beta)$ is the prior distribution of the parameters $\theta$, $\psi$ and $\beta$, and the marginal likelihood
[…]
is related to the complete-data likelihood (5) via integration. Equation (8) integrates over all the possible clinical histories of the individuals consistent with the observed data. No general closed-form expression for this integral exists in the semi-Markov case (Hsieh et al., 2002), and numerical integration over this high-dimensional latent space is intractable.
To overcome this difficult integration step, data-augmented MCMC takes advantage of the tractable form  (5). The observed data $\mathbf{Y}$ are augmented with the latent variables $\mathbf{Z}$ such that $L(\theta,\psi;\mathbf{Z})$ and $f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})$ have closed-form expressions. We can then use Markov chain Monte Carlo to sample from the joint posterior distribution
$\pi(\theta,\psi,\beta,\mathbf{Z}\mid\mathbf{Y})\propto f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})L(\theta,\psi;\mathbf{Z})\pi(\theta,\psi,\beta).$ (9)
Given $M$ MCMC draws $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)\right\}_{m=1}^{M}$, marginalizing out latent variables a posteriori becomes trivial. The values $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)\right\}_{m=1}^{M}$ form an empirical approximation to the original distribution of interest $\pi\left(\theta,\psi,\beta\mid\mathbf{Y}\right)$. In other words, simply ignoring the latent variables in each posterior sample yields a valid marginal posterior of the parameters of interest. Similarly, $\left\{\mathbf{Z}^{(m)}\right\}_{m=1}^{M}$ form a Monte Carlo approximation to $\pi\left(\mathbf{Z}\mid\mathbf{Y}\right)$.
An advantage of the Bayesian formulation is that inference is not limited to $\theta$, $\psi$, $\beta$ and $\mathbf{Z}$. The posterior distribution $\pi\left(g(.)\mid\mathbf{Y}\right)$ of any functional $g(\theta,\psi,\beta,\mathbf{Z})$ is approximated by the transformed sample $\left\{g^{(m)}\right\}_{m=1}^{M}$, with $g^{(m)}=g\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)$. In Section 6, we will be interested in $g$ being the overdiagnosis rate.
In our model, the latent variables $\mathbf{Z}$ consist of the unknown transition ages into $P$ and the indolent status of each individual. We write $\mathbf{Z}=(\mathbf{Z}_{1},\dots,\mathbf{Z}_{n})$, where $\mathbf{Z}_{i}=\left(Z^{HP}_{i},Z^{I}_{i}\right)$ are the latent variables of individual $i$. $Z^{HP}_{i}\in(t_{0},c_{i}]\cup\varnothing$ denotes the pre-clinical onset age and $Z^{I}_{i}\in\{0,1\}$ the indolent status. Here, $\varnothing$ indicates the event that the corresponding transition age is larger than $c_{i}$.
Upon introducing these latent variables, we have closed-form expressions for $L(\theta,\psi;\mathbf{Z})$ and $f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})$. First, by independence, we have
$L(\theta,\psi;\mathbf{Z})=\prod_{i=1}^{n}L_{i}(\theta,\psi;\mathbf{Z}_{i})=\prod_{i=1}^{n}\tilde{L}_{i}(\theta,\psi;\mathbf{Z}_{i})/N_{i}(\theta,\psi,e_{i}),$ (10)
where
$\tilde{L}_{i}(\theta,\psi;\mathbf{Z}_{i})\propto f(\mathbf{Z}_{i}\mid\theta,\psi)=\begin{cases}S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right),&\quad Z^{HP}_{i}=\varnothing\\ f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)f_{\textsf{I}}(Z^{I}_{i};\psi),&\quad Z^{HP}_{i}\leq c_{i}\end{cases},$ (11)
and
$N_{i}(\theta,\psi,e_{i})$ $=\Pr(\tau^{PC}_{i}>e_{i}\mid\theta,\psi)$
$=\Pr(I_{i}=1)+\Pr\left(I_{i}=0,\tau^{HP}_{i}>e_{i}\right)+\Pr\left(I_{i}=0,\tau^{HP}_{i}<e_{i}<\tau^{PC}_{i}\right)$
$=\Pr(I_{i}=1)+\Pr(I_{i}=0)\left[\Pr\left(\tau^{HP}_{i}>e_{i}\mid I_{i}=0\right)+\Pr\left(\tau^{HP}_{i}<e_{i},e_{i}<\tau^{PC}_{i}\mid I_{i}=0\right)\right]$
[…]
scale the contribution from person $i$ by the probability that they did not develop clinical cancer before entering the study at age $e_{i}$. The inclusion of the factor (3.3) in the likelihood (10) comes from the fact that data from cohort analyses are left-truncated: study participants that developed clinical cancer before the study are excluded. Therefore, we condition on the fact that the individuals in the sample have not developed clinical cancer before entering the study. As long as $f_{\textsf{H}}$ and $S_{\textsf{prog}}$ are continuous and bounded, the univariate integral in (3.3) is easy to approximate numerically using a quadrature rule (e.g., the function integrate in base R).
Second, by independence, we have
$f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})=\prod_{i=1}^{n}f(Y_{i}\mid\mathbf{Z}_{i},\theta,\psi,\beta),$ (13)
where
$f(Y_{i}\mid\mathbf{Z}_{i},\theta,\psi,\beta)=f_{t}\left(t^{PC}_{i}\mid\mathbf{Z}_{i},\theta\right)f_{O}\left(O^{i}_{1},\dots,O^{i}_{n_{i}}\mid\mathbf{Z}_{i},\beta\right);$ (14)
the distribution of the right-censored clinical onset age is
$f_{t}\left(t^{PC}_{i}\mid\mathbf{Z}_{i},\theta\right)=\begin{cases}\delta_{c_{i}}\left(t^{PC}_{i}\right),&\quad Z^{HP}_{i}=\varnothing\\ S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}\delta_{c_{i}}\left(t^{PC}_{i}\right)+f_{\textsf{prog}}\left(t^{PC}_{i}-Z^{HP}_{i};\theta\right)\delta_{\tau^{PC}_{i}}\left(t^{PC}_{i}\right),&\quad Z^{HP}_{i}\leq c_{i}\end{cases},$ (15)
and the distribution of the screen outcomes is
$f_{O}\left(O^{i}_{1},\dots,O^{i}_{n_{i}}\mid\mathbf{Z}_{i},\beta\right)=\prod_{j=1}^{n_{i}}\left(p^{i}_{j}\right)^{O^{i}_{j}}\left(1-p^{i}_{j}\right)^{1-O^{i}_{j}}=\beta^{m_{i}^{+}}(1-\beta)^{m_{i}^{-}},$ (16)
where $m_{i}^{+}=\sum_{j=1}^{n_{i}}O^{i}_{j}$ denotes the number positive screens of person $i$, and $m_{i}^{-}=\sum_{j:Z^{HP}_{i}\leq t^{i}_{j}}\left(1-O^{i}_{j}\right)$ their number of false negative screens.
3.3.1 Prior specification
While the sampling methodology applies to any prior formulation $\pi(\theta,\psi,\beta)$, we assume for convenience here that the parameters are independent a priori
$\pi(\theta,\psi,\beta)=\pi(\theta)\pi(\psi)\pi(\beta),$ (17)
Moreover, we specify the prior distribution of $\beta$ to be $\pi(\beta)=\textsf{Be}(a_{\beta},b_{\beta})$ and that of $\psi$ to be $\pi(\beta)=\textsf{Be}(a_{\beta},b_{\beta})$, where $\textsf{Be}(a,b)$ denotes to the beta distribution with density proportional to $x^{a-1}(1-x)^{b-1}$. The choice of a beta prior distribution for $\beta$ is motivated by conjugacy. Theorem 1 holds for any choice of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$.
Theorem 1 (Conjugacy).
If $\beta$ is independent of the other parameters a priori, then the beta distribution is conjugate to the conditional likelihood in (9) for $\beta$. In particular, under the prior distribution
$\pi(\theta,\psi,\beta)=\pi(\theta,\psi)\pi(\beta),\quad\pi(\beta)=\textsf{Be}(a,b),$ (18)
the full conditional distribution of $\beta$ is
$\pi(\beta\mid\theta,\psi,\mathbf{Z},\mathbf{Y})=\textsf{Be}\left(a+m^{+},b+m^{-}\right),$ (19)
with $m^{+}=\sum_{i=1}^{n}m_{i}^{+}$ and $m^{-}=\sum_{i=1}^{n}m_{i}^{-}$ the total number of true positive and false negative screens in the sample, respectively.
Proof.
The complete proof appears in Appendix A. ∎
4 Markov chain Monte Carlo scheme
We propose a data-augmented Markov chain Monte Carlo sampler that alternates between a Gibbs update for $\beta$, Metropolis-Hastings steps for the elements of $\theta$, and for $Z^{HP}_{1},\dots,Z^{HP}_{n}$, and a block Metropolis-Hastings update of $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$. Updating $\psi$ and $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)$ jointly is advantageous because we find these quantities are highly correlated a posteriori. This sampling scheme is summarized as a pseudo-algorithm in Algorithm 1. We note that methodologically, the elements of $\theta$ could also be jointly updated in a block Metropolis-Hastings step which may lead to further efficiency gains, though we do not explore this here.
Algorithm 1 Data-augmented Markov chain Monte Carlo
a number of iterations $M$, initial values $(\theta^{(0)},\psi^{(0)},\beta^{(0)})$, and observed data $\mathbf{Y}$
$\mathbf{Z}^{(0)}\leftarrow\text{initialize}\left(\theta^{(0)},\psi^{(0)},\beta^{(0)}\right)$
for $m=1,\dots,M$ do
     $\beta^{(m)}\sim\pi\left(\beta\mid\cdot\right)$  [Gibbs update, c.f. Theorem 1]
     for $j=1,\dots,|\theta|$ do
         $\theta_{j}^{(m)}\leftarrow\text{Metropolis-Hasting}\left(\theta_{j};.\right)$  [update the $j$th element of $\theta$]
     end for
     for $i=1,\dots,n$ do
         $\left(Z^{HP}_{i}\right)^{(m)}\leftarrow\text{Metropolis-Hasting}\left(Z^{HP}_{i};.\right)$
     end for
     $\psi^{(m)},\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{(m)}\leftarrow\text{Metropolis-Hasting}\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n};.\right)$
end for
return $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)\right\}_{m=1}^{M}$
Next, we briefly review Gibbs and Metropolis-Hastings updates (Brooks et al., 2011). The Gibbs step for $\beta$ proceeds by drawing a new value for this parameter from its full conditional distribution (19) given the most recent values of $\mathbf{Z}$. For the Metropolis-Hasting steps, consider the random vector $(\zeta,\eta)$ where $\zeta$ and $\eta$ are random vectors and from which we wish to update $\zeta$. Given current values $(\tilde{\zeta},\tilde{\eta})$, a candidate value $\zeta^{\star}$ is drawn according to the proposal density $q(\zeta\mid\tilde{\zeta},\tilde{\eta})$, and is accepted with probability
$\alpha\left(\tilde{\zeta},\tilde{\eta}\rightarrow\zeta^{\star},\tilde{\eta}\right)=\min\left\{1,\dfrac{\pi\left(\zeta^{\star},\tilde{\eta}\right)q\left(\tilde{\zeta}\mid\zeta^{\star},\tilde{\eta}\right)}{\pi\left(\zeta^{(m)},\tilde{\eta}\right)q\left(\zeta^{\star}\mid\tilde{\zeta},\tilde{\eta}\right)}\right\}=\min\left\{1,\dfrac{\pi\left(\zeta^{\star}\mid\tilde{\eta}\right)q\left(\tilde{\zeta}\mid\zeta^{\star},\tilde{\eta}\right)}{\pi\left(\zeta^{(m)}\mid\tilde{\eta}\right)q\left(\zeta^{\star}\mid\tilde{\zeta},\tilde{\eta}\right)}\right\},$ (20)
where $\pi(\eta,\zeta)$ is the target distribution; otherwise, the current value is kept. The proposal distributions for $Z^{HP}_{i}$ and $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$ are described in Sections 4.1 are 4.2 respectively. Naturally, the proposal distributions for $\theta$ depend on the choice of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$.
4.1 Metropolis-Hastings update for $Z^{HP}_{i}$
The efficiency of the Metropolis-Hastings steps for $Z^{HP}_{1},\dots,Z^{HP}_{n}$ is central to the overall efficiency of our data-augmented Markov chain Monte Carlo sampler. We construct a proposal distribution for $Z^{HP}_{i}$ that is independent of the current value of this latent variable; a higher acceptance rate, therefore, results in better mixing. To this end, we design a proposal distribution that is close to the full conditional $\pi\left(Z^{HP}_{i}\mid\cdot\right)$. The support of $\pi\left(Z^{HP}_{i}\mid\cdot\right)$ is $(t_{0},c_{i}]\cup\varnothing$ for individuals in the censored group and $(t_{0},c_{i}]$ for those in the screen-detected and interval-detected groups. We therefore consider a proposal distribution with a positive density on the interval $(t_{0},c_{i}]$ and a point mass at $\varnothing$ for the censored group. The exact form of the proposal distribution depends on the group of the individual, described below.
4.1.1 Censored group
For an individual $i$ in the censored group, the observed data consist of $n_{i}$ negative screens and $t^{PC}_{i}=c_{i}$. The support of the full conditional distribution of $Z^{HP}_{i}$ is $(t_{0},c_{i}]\cup\varnothing$, and from Equations (9), (11) and (14), (15), its density is
$\pi\left(Z^{HP}_{i}\mid\cdot\right)\propto L_{i}(\theta,\psi;Z^{HP}_{i},Z^{I}_{i})f_{t}\left(t^{PC}_{i}\mid Z^{HP}_{i},Z^{I}_{i},\theta\right)f_{O}\left(O^{i}_{1},\dots,O^{i}_{n_{i}}\mid Z^{HP}_{i},\beta\right)$
$\quad\propto\left[\delta_{\varnothing}\left(Z^{HP}_{i}\right)S_{\textsf{H}}(c_{i}-t_{0})+\delta_{(t_{0},c_{i})}\left(Z^{HP}_{i}\right)f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}\right]\beta^{m_{i}^{+}}(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}$
$\quad=\delta_{\varnothing}\left(Z^{HP}_{i}\right)S_{\textsf{H}}(c_{i}-t_{0})+\delta_{(t_{0},c_{i})}\left(Z^{HP}_{i}\right)f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}.$ (21)
Here $m_{i}^{+}=0$ because person $i$ belongs to the censored group, the notation $m_{i}^{-}\left(Z^{HP}_{i}\right)$ emphasizes that the number of false negative screens is a function of $Z^{HP}_{i}$, and the equality holds because $m_{i}^{-}\left(\varnothing\right)=0$. The full conditional (4.1.1) has a point mass at $\varnothing$ and a positive density on $(t_{0},c_{i}]$. The integer-valued quantity $m^{-}\left(Z^{HP}_{i}\right)$ induces discontinuities, as shown in Figure 3(a).
We construct a proposal distribution for the Metropolis-Hastings step that mimics the full conditional  (4.1.1). First, the proposal has a point mass component at $\varnothing$ with weight proportional to $S_{\textsf{H}}(c_{i}-t_{0})$. The density component is constructed as a mixture distribution with a discontinuous density proportional to $f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}$. The expression is designed to be equal to the full conditional distribution whenever the pre-clinical cancer is indolent ($Z^{I}_{i}=1$), and otherwise only differs by the factor $S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)$ for the case of progressive pre-clinical cancer. The mixture components are truncated Weibull distributions restricted to non-overlapping supports defined by the inter-screen intervals. Over the $k$th inter-screen interval $(t^{i}_{k-1},t^{i}_{k}]$, the truncated Weibull density of the corresponding mixture component is proportional to $f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)$, and its mixture weight is proportional to
[…]
We choose to construct a mixture distribution whose continuous component has a density proportional to $f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}$ instead of $f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}$ because the extra term $S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}$ makes the mixture components and weights intractable.
Putting everything together gives the mixed proposal distribution
$q\left(Z^{HP}_{i}\mid\cdot\right)$ $=p^{i}_{0}\delta_{\varnothing}\left(Z^{HP}_{i}\right)+\sum_{k=1}^{n_{i}+1}p^{i}_{k}\textsf{Trunc-F}_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta,t^{i}_{k-1}-t_{0},t^{i}_{k}-t_{0}\right)$ (22)
[…]
$p^{i}_{0}$ $\propto S_{\textsf{H}}(c_{i}-t_{0}),$ (23)
$p^{i}_{k}$ $\propto(1-\beta)^{n_{i}+1-k}\left[F_{\textsf{H}}\left(t^{i}_{k}-t_{0};\theta\right)-F_{\textsf{H}}\left(t^{i}_{k-1}-t_{0};\theta\right)\right],\quad k=1,\dots,n_{i}+1.$ (24)
Here $F_{\textsf{H}}$ is the cumulative distribution function of $\textsf{F}_{\textsf{H}}$. Because $p^{i}_{k}>0,\forall k$, Equations (22)-(24) define a mixed distribution with a strictly positive density on its support $(t_{0},c_{i}]\cup\varnothing$.
We sample from this distribution in a two-step procedure. First, we draw the index $k\in\{0,\dots,n_{i}+1\}$ from the categorical distribution with weights $p_{0},\dots,p_{n_{i}+1}$. The index $k=0$ corresponds to the event that the $P$ transition occurs after the censoring time, $k=1$ that it occurs occurs between $t_{0}$ and the first screen, the index $k=2,\dots,n_{i}$ that it occurs between the $(k-1)$th and the $k$th screen, and $k=n_{i}+1$ that it occurs between the last screen and the censoring age. Second, we sample $Z^{HP}_{i}$ conditionally on $k$ as follows
$Z^{HP}_{i}\mid k\sim\begin{cases}\delta_{\varnothing}\left(Z^{HP}_{i}\right),\quad&k=0\\ \textsf{Trunc-F}_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta,t^{i}_{k-1}-t_{0},t^{i}_{k}-t_{0}\right),\quad&k=1,\dots,n_{i}+1\end{cases},$ (25)
This produces a proposal from (22), which is then accepted or rejected according to Equation (20) in the Metropolis-Hastings step.
4.1.2 Screen-detected cancer group
The observed data for individual $i$ in the screen-detected group consist of negative screens followed by a positive screen at $t^{i}_{n_{i}}=c_{i}$. The support of the full conditional distribution of $Z^{HP}_{i}$ is $(t_{0},c_{i}]$, and its density is
$\pi\left(Z^{HP}_{i}\mid\cdot\right)\propto f_{H}\left(Z^{HP}_{i}-t_{0};\theta\right)S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}\beta^{m^{+}}(1-\beta)^{m^{-}\left(Z^{HP}_{i}\right)},$ (26)
where $\beta^{m^{+}}=\beta$ is a constant which we can drop by proportionality.
Following the same reasoning as in Section 4.1.1, we construct a proposal distribution consisting of a mixture of non-overlapping truncated Weibull distribution mimicking the full conditional (26). The proposal distribution is
$q\left(Z^{HP}_{i}\mid\cdot\right)=\sum_{k=1}^{n_{i}}p^{i}_{k}\textsf{Trunc-F}_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta,t^{i}_{k-1}-t_{0},t^{i}_{k}-t_{0}\right),$ (27)
where $t^{i}_{0}=t_{0}$, and the mixture weights are
$p^{i}_{k}$ $\propto(1-\beta)^{n_{i}-k}\left[F_{\textsf{H}}\left(t^{i}_{k}-t_{0};\theta\right)-F_{\textsf{H}}\left(t^{i}_{k-1}-t_{0};\theta\right)\right],\quad k=1,\dots,n_{i}.$ (28)
Equations (27)-(28) define a mixture distribution with a positive density on $(t_{0},c_{i}]$ from which we sample following the two-step procedure described in Section 4.1.1.
4.1.3 Interval cancer group
In the interval cancer group, the observed data for individual $i$ consist of a series of negative screens followed by a clinical diagnosis at time $t^{PC}_{i}=\tau^{PC}_{i}$. Moreover, we have $Z^{I}_{i}=0$ because the individual’s pre-clinical cancer was necessarily progressive. The support of the full conditional distribution of $Z^{HP}_{i}$ is $(t_{0},c_{i}]$, and its density is
$\pi\left(Z^{HP}_{i}\mid\cdot\right)$ $\propto f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)f_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)(1-\beta)^{m^{-}\left(Z^{HP}_{i}\right)}.$ (29)
As a proposal distribution, we employ the mixture distribution with density
$q\left(Z^{HP}_{i}\mid\cdot\right)=\sum_{k=1}^{n_{i}+1}p^{i}_{k}\textsf{Trunc-F}_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta,t^{i}_{k-1}-t_{0},t^{i}_{k}-t_{0}\right),$ (30)
where $t^{i}_{0}=t_{0}$ and $t^{i}_{n+1}=\tau^{PC}_{i}$, and the mixture weights are
$p^{i}_{k}$ $\propto(1-\beta)^{n_{i}+1-k}\left[F_{H}\left(t^{i}_{k}-t_{0}\right)-F_{H}\left(t^{i}_{k-1}-t_{0}\right)\right],\quad k=1,\dots,n_{i}+1.$ (31)
Equations (30)-(31) define a mixture distribution with a positive density on $\left(t_{0},\tau^{PC}_{i}\right]$ from which we sample following the two-step procedure described in Section 4.1.1.
As illustrated in Figure 3, the three proposal distributions (22), (27) and (30) all closely match their corresponding full conditional distributions, resulting in a high acceptance rate in the Metropolis-Hastings steps (see also the simulations and case study in Sections 5-6). In the case of a progressive pre-clinical cancer, the proposal distribution—which ignores the term $S$—is close to the full conditional distribution, and in the case of an indolent pre-clinical cancer, the proposal distribution matches exactly the full conditional distribution (see Figures 3(a)-3(b)). Moreover, the proposals do not depend on $Z^{I}_{i}$, a fact which will be useful in Section 4.4.
(a) Censored group.
(b) Screen-detected group.
(c) Interval-detected group.
Figure 3: Comparison of the proposal and target densities in the Metropolis-Hastings step for the latent variable $Z^{HP}_{i}$. Solid line: proposal distribution $q\left(Z^{HP}_{i}\mid\cdot\right)$; dashed line: conditional distribution for progressive cancer $\pi\left(Z^{HP}_{i}\mid Z^{I}_{i}=0,.\right)$; long-dashed line: conditional distribution for indolent cancer $\pi\left(Z^{HP}_{i}\mid Z^{I}_{i}=1,.\right)$. In Panels (a) and (b), the proposal distribution matches exactly the conditional distribution for indolent cancers.
4.2 Joint Metropolis-Hastings update for $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$
We update $\psi$ and the vector $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)$ jointly because these two objects are highly correlated a posteriori due to the observed data providing little information about the latent indolent status of censored and screen-detected individuals. At iteration $m=1,\dots,M$ of the sampler, we use the two-step proposal distribution
$q\left(\psi^{\star},Z^{I}_{1},\dots,Z^{I}_{n}\mid\psi^{(m)},.\right)=q_{\psi}\left(\psi^{\star}\mid\psi^{(m)}\right)q_{I}\left(\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{\star}\mid\psi^{\star},.\right),$ (32)
which first samples a new value $\psi^{\star}$ and then samples a new vector $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{\star}$ conditionally on $\psi^{\star}$.
The proposal distribution $q_{\psi}\left(\psi^{\star}\mid\psi^{(m)}\right)$ corresponds to a Gaussian random walk with reflection at 00 and $1$, ensuring that $\psi^{\star}\in[0,1]$. Simple algebra shows that the proposal distribution $q_{\psi}$ is symmetric, and its variance $\epsilon$ is a tuning parameter. If $\epsilon$ is too small, the Markov chain will take small steps, inefficiently exploring the space of $\psi$, while if $\epsilon$ is too large, most proposed values will be rejected, resulting in a Markov chain that is often stuck at the same location. In our implementation of the sampler, we tune the value of $\epsilon$ to reach a pre-specified acceptance rate during an warmup phase using the adaptive procedure described in (Hoffman et al., 2014), and then fix $\epsilon$ for the remaining iterations. Details of the adaptive procedure are provided in Appendix E.
Next, $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{\star}$ is sampled conditionally on $\psi^{\star}$ by drawing each $Z^{I}_{i}$ independently from its full conditional distribution given $\psi^{\star}$. If a person $i$ is diagnosed with an interval cancer, we set $Z^{I}_{i}=0$ as their cancer is necessarily progressive; otherwise
$\pi(Z^{I}_{i}\mid\psi^{\star},.)=\textsf{Ber}(q_{i}),$ (33)
where
$q_{i}=\Pr\left(Z^{I}_{i}=1\mid\psi^{\star},\cdot\right)=\frac{L_{i}\left(Z^{I}_{i}=1;\psi^{\star},\cdot\right)}{L_{i}\left(Z^{I}_{i}=0;\psi^{\star},\cdot\right)+L_{i}\left(Z^{I}_{i}=1;\psi^{\star},\cdot\right)},$ (34)
with
$L_{i}(Z^{I}_{i};\psi^{\star},\cdot)=\begin{cases}\textsf{Ber}(Z^{I}_{i};\psi^{\star}),\quad&Z^{HP}_{i}=\varnothing\\ \textsf{Ber}(Z^{I}_{i};\psi^{\star})S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}},\quad&Z^{HP}_{i}\leq c_{i}\end{cases},$ (35)
the likelihood of $Z^{I}_{i}$.
4.3 Metropolis-Hastings update for $\theta$
We update the elements of $\theta$ in separate univariate Metropolis-Hastings steps using a Gaussian random walk proposal. Further complete details of these steps are described in Appendix B.
4.4 Initialization of the sampler
To initialize the sampling algorithm, one only needs to specify the initial values of $(\theta,\psi,\beta)$. Given some initial values $\left(\theta^{(0)},\psi^{(0)},\beta^{(0)}\right)$, we generate the initial values of the latent variables using the proposal distributions of the Metropolis-Hastings steps. Specifically, the initial values ${Z^{HP}_{1}}^{(0)},\dots,{Z^{HP}_{n}}^{(0)}$ are generated independently using the distributions (22), (27) and (30) based on the group—censored, screen-detected cancer or interval cancer—of each individual, and $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{(0)}$ are generated from (33) independently.
4.5 Theoretical guarantees
Theorem 2 shows that the Markov chain underlying the data-augmented Markov chain Monte Carlo sampler is ergodic. The ergodic theorem, therefore, holds for any $\pi$-integrable function $h$, and the estimator $\bar{h}_{m}=\frac{1}{m}\sum_{i=1}^{m}h\left(\theta^{(i)},\beta^{(i)},\mathbf{Z}^{(i)}\right)$ is asymptotically consistent for $E_{\pi}h$ under any initial distribution (Billingsley, 2017).
Theorem 2 (Ergodicity).
The Markov chain $\left\{\left(\theta^{(m)},\theta^{(m)},z^{(m)}\right)\right\}_{m=1}^{M}$ underlying the data-augmented Markov chain Monte Algorithm algorithm described in this article is ergodic.
Proof.
We follow the proof of Morsomme and Xu (2022). By construction, the distribution $\pi(\theta,\psi,\beta,\mathbf{Z}\mid\mathbf{Y})$ is invariant for the kernels $P_{\beta}$, $P_{\theta}$, $P_{Z^{HP}_{i}}$ (for $i=1,\dots,n$) and $P_{\psi,Z^{I}}$ that respectively update $\beta$, $\theta$, $Z^{HP}_{i}$ and $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$. The distribution $\pi(\theta,\psi,\beta,\mathbf{Z}\mid\mathbf{Y})$ is, therefore, also invariant for the composite kernel $P=P_{\beta}P_{\theta}P_{Z^{HP}_{1}}\dots P_{Z^{HP}_{n}}P_{\psi,Z^{I}}$ (Tierney, 1994). Moreover, as each kernel is strictly positive, $P$ is also strictly positive. As a result, the chain is $\pi$-irreducible and aperiodic. The chain is, therefore, positive Harris recurrent and thus ergodic. ∎
4.6 Model comparison
Mis-specifying a model of cancer natural history can lead to biased inference (Cheung et al., 2022). To identify specifications for $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ that fit the data well, we compare the predictive fit of models with different choices for these distributions. We accomplish this by extending approximate leave-one-out cross-validation (ALOOCV) Gelfand et al. (1992); Gelfand (1995); Vehtari et al. (2017) to the class of latent variable models. While exact cross-validation requires fitting the model multiple times, approximate cross-validation can be easily computed from the posterior draws of a MCMC sampler. The procedure is straight-forward to implement, and enables a hypothesis test for the difference in predictive fit between models.
We define the leave-one-out out-of-sample predictive fit as
$\textsf{pf}=\sum_{i=1}^{n}\log(f(y_{i}\mid\mathbf{Y}_{-i})),$ (36)
where $\mathbf{Y}_{-i}=(Y_{1},\dots,Y_{i-1},Y_{i+1},\dots,Y_{n})$ denotes the data set without the $i$th subject and
[…]
is the predictive density for subject $i$ given the data of the other subjects, and prefer model with larger predictive fit.
We approximate Eq (37) using the importance sampling estimator
$\hat{f}(y_{i}\mid\mathbf{Y}_{-i})=\frac{\sum_{m=1}^{M}r^{(m)}_{i}f(y_{i}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)})}{\sum_{m=1}^{M}r^{(m)}_{i}},$ (38)
with importance weight $r^{(m)}_{i}=\frac{\pi(\theta,\psi,\beta\mid\mathbf{Y}_{-i})}{g\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)}$ for some proposal distribution $g$.
Vehtari et al. (2017) suggest to choose
$g(\theta,\psi,\beta)=\pi(\theta,\psi,\beta\mid\mathbf{Y})$ (39)
for two reasons. First, for large samples $\pi(\theta,\psi,\beta\mid\mathbf{Y})$ tends to be similar to $\pi(\theta,\psi,\beta\mid\mathbf{Y}_{-i})$, making it a suitable proposal distribution for importance sampling. Second, by construction, once the Markov chain has reached stationarity the draws $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)\right\}_{m}$ come from $\pi(\theta,\psi,\beta\mid\mathbf{Y})$, making the proposal (39) straightforward to implement. This choice of proposal results in the importance weight
$r^{(m)}_{i}$ $=\frac{\pi\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\mid\mathbf{Y}_{-i}\right)}{\pi\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\mid\mathbf{Y}\right)}$
$\propto\frac{f\left(\mathbf{Y}_{-i}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)}{f\left(\mathbf{Y}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)}$
$=\frac{1}{f\left(y_{i}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)},$
giving the estimator
$\hat{f}(y_{i}\mid\mathbf{Y}_{-i})=\left(\frac{1}{M}\sum_{m=1}^{M}\frac{1}{f\left(y_{i}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)}\right)^{-1},$ (40)
where $\left\{\left(\theta^{(1)},\psi^{(1)},\beta^{(1)}\right)\,\dots,\left(\theta^{(M)},\psi^{(M)},\beta^{(M)}\right)\right\}$ correspond to the posterior draws of the MCMC sampler. Finally, we estimate the out-of-sample predictive fit (36) with
$\widehat{\textsf{pf}}=\sum_{i=1}^{n}\log(\hat{f}(y_{i}\mid\mathbf{Y}_{-i}))=-\sum_{i=1}^{n}\log\left(\frac{1}{M}\sum_{m=1}^{M}\frac{1}{f\left(y_{i}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)}\right).$ (41)
4.6.1 Estimating $f(y_{i}\mid\theta,\psi,\beta)$
For our model, the likelihood $f\left(y_{i}\mid\theta,\psi,\beta\right)$ does not admit a closed-form expression. We use importance sampling to estimate this quantity, taking advantage of the expression
[…]
Equation (4.6.1) suggests the importance sampling estimate
$\hat{f}\left(y_{i}\mid\theta,\psi,\beta\right)$ $=\frac{1}{J}\sum_{j=1}^{J}f\left(y_{i}\mid\theta,\psi,\beta,Z_{i}^{j}\right)w^{j},$
with importance weights $w^{j}=\frac{f\left(Z_{i}^{j}\mid\theta,\psi,\beta\right)}{q\left(Z_{i}^{j}\right)}$ for some proposal distribution $q$.
We choose $q$ to correspond to the Metropolis-Hastings proposal distributions for the latent variables described in Section 4.1-4.2,
$q\left(Z_{i}\right)=q_{\textsf{HP}}\left(Z^{HP}_{i}\mid\cdot\right)q_{\textsf{I}}\left(Z^{I}_{i}\mid Z^{HP}_{i},\cdot\right),$ (43)
with $q_{\textsf{HP}}$ corresponding to Equation 22, 27 or 30 based on the group of indidual $i$, and $q_{\textsf{I}}$ corresponding to Equation 33. We sample $Z_{i}\sim q$ sequentially, first sampling $Z^{HP}_{i}\sim q_{\textsf{HP}}$ and then sampling $Z^{I}_{i}\mid Z^{HP}_{i}\sim q_{\textsf{I}}$.
Substituting $\hat{f}\left(y_{i}\mid\theta,\psi,\beta\right)$ for $f\left(y_{i}\mid\theta,\psi,\beta\right)$ in (41) gives the predictive fit estimate
$\widehat{\textsf{pf}}=\sum_{i=1}^{n}\log(\hat{f}(y_{i}\mid\mathbf{Y}_{-i}))=-\sum_{i=1}^{n}\log\left(\frac{1}{M}\sum_{m=1}^{M}\frac{1}{\frac{1}{J}\sum_{j=1}^{J}f\left(y_{i}\mid\theta^{(m)},\psi^{(m)},\beta^{(m)},Z_{i}^{(j)}\right)w^{j}}\right).$ (44)
Appendix C describes a $t$-test of the hypothesis that two models have the same predictive fit (36).
4.7 Model implementation
[…]
We implement the MCMC sampler in the programming language C++ for efficiency. The sampler is fast, scaling to cohort studies with tens of thousands of individuals. For convenience, the C++ implementation is made available in R via the R package baclava with the help the package RCPP (Eddelbuettel and François, 2011). In the MCMC runs in Sections 5-6, we keep $10^{3}$ thinned draws post warm-up phase due to memory constraints.
5 Performance on simulated data
To examine the mixing properties of our sampler, we run multiple MCMC samplers initialized at over-dispersed values on a simulated dataset. We then use the potential scale reduction factor (Gelman and Rubin, 1992; Brooks and Gelman, 1998), the effective sample size and traceplors to assess the algorithm’s mixing properties.
For the purposes of a synthetic dataset, we set the starting age for non-zero onset hazard to $t_{0}=30$ years, and assume a linear increase in both the onset and progression hazards ($\alpha_{H}=\alpha_{prog}=2$). We further set the Weibull rates to $\lambda_{\textsf{H}}=6.5\times 10^{-5}$ and $\lambda_{\textsf{prog}}=3.14\times 10^{-2}$, resulting in a cumulative risk of pre-clinical onset by age $80$ of $15\%$, and a mean sojourn time of $\mu_{\textsf{prog}}=5$ years. Finally, we set the indolent probability to $\psi=0.1$ and the screen sensitivity to $\beta=0.85$. We simulate the natural histories of cancer among $n=40,000$ individuals (see Appendix D for the distributions of screening and censoring ages).
We employ the independent prior distributions $\pi(\lambda_{\textsf{H}},\lambda_{\textsf{prog}},\psi,\beta)=\pi(\lambda_{\textsf{H}})\pi(\lambda_{\textsf{prog}})\pi(\psi)\pi(\beta)$ with
$\pi(\lambda_{\textsf{H}})=\textsf{Ga}(1,0.01),\quad\pi(\lambda_{\textsf{prog}})=\textsf{Ga}(1,0.01),\quad\pi(\psi)=\textsf{Be}(1,1),\quad\pi(\beta)=\textsf{Be}(38.5,5.8)$
We choose these prior distributions for their flexibility and, in the case of $\pi(\beta)$, for its conjugacy (Theorem 1). The prior distribution on $\lambda_{\textsf{H}}$, $\lambda_{\textsf{prog}}$ and $\psi$ are weakly informative, and that of $\beta$ places $95\%$ of its mass on the interval $(0.76,0.95)$, reflecting prior empirical data on the performance of mammography screening of clinicians about that parameter (Lehman et al., 2017).
Next, we apply our MCMC sampler to the synthetic dataset and estimate the values of the parameters $\lambda_{\textsf{H}}$, $\lambda_{\textsf{prog}}$, $\psi$ and $\beta$. We independently run $20$ MCMC samplers initialized at over-dispersed values of these parameters (the latent variables need not be initialized, see Section 4.4). The initial values $\lambda_{\textsf{H}}^{(0,m)}$, $\lambda_{\textsf{prog}}^{(0,m)}$, $\psi^{(0,m)}$ and $\beta^{(0,m)}$ of the $m$th sampler $(m=1\dots,20)$ are sampled from a wide interval around the true values, mutually independently,
$\lambda_{\textsf{H}}^{(0,m)}\sim\textsf{Unif}(2\times 10^{-5},2\times 10^{-4}),\quad\lambda_{\textsf{prog}}^{(0,m)}\sim\textsf{Unif}(10^{-2},10^{-1}),\quad\psi^{(0,m)}\sim\textsf{Unif}(0,1),\quad\beta^{(0,m)}\sim\textsf{Unif}(0.7,0.95).$
For each chain, we set the warm-up period to $10^{4}$ iterations and run the sampler for an additional $2\times 10^{4}$ iterations. Table 1 shows that the potential scale reduction factor, computed on the post warm-up thinned draws with the R package coda (Plummer et al., 2006), is smaller than $1.005$ for all parameters, indicating that the chains—which were initialized at over-dispersed values—converged during the warmup phase and mixed adequately once they reached stationarity (Gelman and Rubin, 1992). This is a remarkable achievement considering the high dimensional latent space with 80,000 latent variables.
Parameter Potential scale reduction factor Effective sample size
$\lambda_{\textsf{H}}$ $1.0005$ $11,728$
$\lambda_{\textsf{prog}}$ $1.0027$ $2,731$
$\psi$ $1.002$ $2,568$
$\beta$ $1.0013$ $6,497$
Table 1: Performance of the MCMC sampler on simulated data.
In addition to exploring the parameter space efficiently, the sampler also explores the space of the latent variables swiftly. Figure 4 displays the MCMC draws of the latent variables for an individual whose cancer was screen-detected at age $58$ years, after a series of five consecutive negative screens. The individual had $5$ negative screens at ages . The traceplots of the indolent indicator (Figure 4(a)) and the pre-clinical transition time (Figure 4(b)) indicate that the sampler rapidly explores the latent space. In particular, the frequent switches between the values 00 and $1$ of the binary draws in Figure 4(a) show that the Markov chain is not stuck in any particular configuration of the latent variables for any large number of iterations.
An interesting feature of our data augmentation approach is that we obtain a patient-specific estimate of the onset time of pre-clinical cancer, a quantity that is unobservable in practice. Figure 4(c) illustrates the MCMC approximation of the marginal posterior distribution $\pi\left(\tau^{HP}\mid\mathbf{Y}\right)$ for the above individual with a screen-detected cancer at age 58 years. Because the screening sensitivity is high, it is most likely that the time of onset took place after the last screen at $55$ years, less likely that it did between the screens at ages $53$ and $55$, and extremely unlikely before the screen at age $53$. As illustrated by the discontinuities in this histogram, the screens themselves contribute most of the information available about the unobserved onset age of pre-clinical cancer.
(a) Traceplot of $Z^{I}$.
(b) Traceplot of $Z^{HP}$.
(c) Histogram of $Z^{HP}$.
Figure 4: Markov chain Monte Carlo draws from the first chain ($m=1$) for the latent variables of an individual with a screen-detected cancer at age $58$ and five negative screens at ages $(46,48,51,53,55)$. Panel (a) is a traceplot of the latent indolent indicator over an interval of $5.000$ iterations for clarity. Panel (b) is a traceplot of the latent pre-clinical transition age. Panel (c) shows the marginal posterior distribution of the latent pre-clinical transition age; vertical lines depict the ages at which the screens occur.
6 Application to data from the Breast Cancer Surveillance Consortium
We illustrate the utility of our semi-Markov model and accompanying data-augmented MCMC sampler with an analysis of the screening and cancer diagnosis histories from participants in the Breast Cancer Surveillance Consortium (BCSC), a large population-based mammography screening registry in the US (Ballard-Barbash et al., 1997). The BCSC registries collect mammography results and link them to breast outcomes and vital status through linkage with regional population-based SEER (Surveillance, Epidemiology, and End Results) programs, state tumor registries, and state death records.
We used the analytic cohort previously described in Ryser et al. (2022), including $n=35,986$ women aged $50$ to $74$ years who received their first screening mammogram at a BCSC facility between $2000$ and $2018$ (Table 2). Each BCSC registry and the Statistical Coordinating Center (SCC) have received institutional review board approval for all study procedures, including passive permission (one registry), a waiver of consent (six registries), or both depending on facility (one registry), to enroll participants, link data, and perform analytic studies. All procedures are Health Insurance Portability and Accountability Act (HIPAA) compliant. All registries and the SCC have received a Federal Certificate of Confidentiality and other protections for the identities of women, physicians, and facilities who are subjects of this research.
Following Ryser et al. (2022), we set $t_{0}=30$. We employ the independent prior distributions $\pi(\lambda_{\textsf{H}},\lambda_{\textsf{prog}},\psi,\beta)=\pi(\lambda_{\textsf{H}})\pi(\lambda_{\textsf{prog}})\pi(\psi)\pi(\beta)$ with
$\pi(\lambda_{\textsf{H}})=\textsf{Ga}(1,0.01),\quad\pi(\lambda_{\textsf{prog}})=\textsf{Ga}(a_{\textsf{prog}},b_{\textsf{prog}}),\quad\pi(\psi)=\textsf{Be}(1,1),\quad\pi(\beta)=\textsf{Be}(38.5,5.8),$
where, for a given $\alpha_{\textsf{prog}}$, we choose values for $a_{p}$ and $b_{p}$ that induce a weakly informative distribution on the mean sojourn time $\mu_{\textsf{prog}}$ with $0.9$ mass in the interval $(1,9)$ years, reflecting prior knowledge about breast cancer progression (Figure 5(a)) . Figure 5(b) shows that the prior distributions induced on $\mu_{\textsf{prog}}$ are similar across values of $\alpha_{\textsf{prog}}$.
Characteristic Value
Median (IQR) age at first screen, years $56$ ($52$-$64$)
Median (IQR) number screens $1$ ($1$-$3$)
Median (IQR) inter-screen interval, years $1.39$ ($1.05$-$2.19$)
Number of interval-detected cancers $73$
Number of screen-detected cancers $645$
Number of censored individuals $35,268$
Table 2: Summary of the data from the Breast Cancer Surveillance Consortium (n=35,986). IQR: Inter-quartile range.
$\alpha_{\textsf{prog}}$ $a_{\textsf{prog}}$ $b_{\textsf{prog}}$
$1$ $2.64$ $5.75$
$1.5$ $1.36$ $4.28$
$2$ $0.9$ $3.56$
(a) Parameters of $\textsf{Ga}(a_{\textsf{prog}},b_{\textsf{prog}})$.
(b) Implied prior distribution of $\mu_{\textsf{prog}}$.
Figure 5: Weakly informative prior distribution of $\lambda_{\textsf{prog}}$.
6.1 Model specification
We consider the $21$ models with $\alpha_{\textsf{H}}\in\{1,1.5,2,2.5,3,4,5\}$ and $\alpha_{\textsf{prog}}\in\{1,1.5,2\}$. We fit each model to the BCSC data and compare them via ALOOCV (see Section 4.6). The warm-up period is set to $2\times 10^{4}$ iterations and is followed by an additional $5\times 10^{4}$ iterations, resulting in an effective sample size of at least $10^{3}$ for each parameter in every setting.
Figure 6 displays the predictive fit of the models. The relation between predictive fit and $\alpha_{\textsf{H}}$ follows a inverted-U shape, and, nominally, the best model fit was achieved at $\alpha_{\textsf{H}}=4$ and $\alpha_{\textsf{prog}}=1$. Models with $\alpha_{\textsf{H}}\leq 1.5$ and $2$ of the $3$ models with $\alpha_{\textsf{H}}=2$ have an estimated predictive fit that is significantly different from that of the best model at the significance level $0.05$. This shows that letting the pre-clinical hazard rate increase at least linearly with age significantly improves the predictive fit of the model for the BCSC data set. However, the exact speed of the increase—linear, quadratic, etc—cannot be inferred from the data, probably because the available screens only cover the range $50$-$85$ years, with $97\%$ of them limited to the range $50$-$75$ years.
Figure 6: Predictive fit of the models on the BCSC data. The black dot denotes the best model. The red squares correspond to models whose predictive score is significantly different from that of the best model, based on a two-sample t-test at the significance level $0.05$). The blue triangle correspond to models whose predictive fit is not significantly different from that of the best model.
In contrast, $\alpha_{\textsf{prog}}$ does not have significant impact on the predictive fit. The three lines in Figure 6, which correspond to the three values of this parameter, are extremely close to one another, and the pairwise differences between the predictive fit of models with the same $\alpha_{\textsf{H}}$ did not reach statistical significance at $\alpha=0.05$. This indicates that the data contain little information about this parameter. This is to be expected because the endpoints of the pre-clinical sojourn times of progressive cancer are mostly unobserved. Indeed, the start age is never exactly observed; furthermore, as the screens are imperfect, this value is not even interval censored. In addition, the end age of the sojourn is exactly observed only for the interval-detected group ($0.18\%$ of the sample) and is right-censored for the other individuals.
This observation drives us to motivate the choice biologically. In particular, we choose a value of $\alpha_{\textsf{prog}}$ that implies realistic predictive distributions of sojourn times among progressive cancers. Conditional on $\lambda_{\textsf{prog}}$, the sojourn time of a progressive cancer—denoted by $\sigma_{\textsf{prog}}$—follows the Weibull distribution $\textsf{Wei}\left(\lambda_{\textsf{prog}},\alpha_{\textsf{prog}}\right)$. Its predictive distribution
[…]
is therefore a rate-mixture of Weibull distributions, in which the Weibull rate is integrated out. We approximate the posterior distribution $\pi\left(\lambda_{\textsf{prog}}\mid\mathbf{Y},\alpha_{\textsf{prog}}\right)$ with the empirical distribution of the MCMC draws $\left\{\lambda_{\textsf{prog}}^{(1)},\dots,\lambda_{\textsf{prog}}^{(M)}\right\}$, giving the Rao-Blackwell estimate of the predictive distribution (45) (Robert et al., 2007)
$\hat{\pi}(\sigma_{\textsf{prog}}\mid\mathbf{Y},\alpha_{\textsf{prog}})=\dfrac{1}{M}\sum_{m=1}^{M}\textsf{Wei}\left(\sigma_{\textsf{prog}};\lambda_{\textsf{prog}}^{(m)},\alpha_{\textsf{prog}}\right).$ (46)
Figure 7(a) presents the estimate (46) for the three models with $\alpha_{\textsf{H}}=2.5$ and $\alpha_{\textsf{prog}}\in\{1,1.5,2\}$. Under $\alpha_{\textsf{prog}}=1$, the predictive distribution has a large amount of mass concentrated at 00 and a thick tail to the right. In contrast, under the semi-Markov model with $\alpha_{\textsf{prog}}\in\{1.5,2\}$ the density of the predictive distribution has a thin tail to the right, and no mass concentrated at 00. Figure 7(b) presents the probabilities that the sojourn time of a future individual is less than $0.5$ year and more than $15$ years for the different values of $\alpha_{\textsf{prog}}$. Under $\alpha_{\textsf{prog}}=1$, these probabilities amount to almost $20\%$ of progressive cancers with an extremely short or long pre-clinical sojourn time, while for the models with $\alpha_{\textsf{prog}}=1.5,2$, they amount to $5.6\%$ and $1.3\%$ respectively. We thus opt for $\alpha_{\textsf{prog}}=2$ because it results in a more realistic model of sojourn times and note that the user can use any other value for this parameter deemed suitable.
(a) Predictive density.
$\alpha_{\textsf{prog}}$ $P(\sigma_{\textsf{prog}}<0.5)$ $P(\sigma_{\textsf{prog}}>15)$
$1$ $0.07$ $0.12$
$1.5$ $0.0217$ $0.0343$
$2$ $0.0066$ $0.0069$
(b) Tail behavior of the predictive distributions.
Figure 7: Predictive distribution of pre-clinical sojourn times among progressive cancers for the Breast Cancer Surveillance Consortium dataset ($\alpha_{\textsf{H}}=2.5$).
Among the models with $\alpha_{\textsf{prog}}=2$, we choose the one with $\alpha_{\textsf{H}}=2.5$ because it agrees with the trend in age-specific data from the Surveillance, Epidemiology, and End Results Program and its predictive fit is not significantly different from that of the best model. Figure 8 presents the output of the MCMC sampler for the chosen model. The posterior mean of the indolence probability $\psi$ is $0.057$, with $95\%$ credible interval ($0.002-0.17$), and that of the mean sojourn time of a progressive cancer $\mu_{\textsf{prog}}$ is $5.63$ years ($95\%$ CI: $4.06-7.11$). Finally, the hazard rate for pre-clinical cancer at age $40$ is estimated to be $0.59\times 10^{-3}$ ($95\%$ CI: $0.53\times 10^{-3}-0.65\times 10^{-3}$) and that at age $80$ is estimated to be $6.6\times 10^{-3}$ ($95\%$ CI: $5.9\times 10^{-3}-7.2\times 10^{-3}$), and the probability of pre-clinical onset by age 80 is estimated to be $12.3\%$ ($95\%$ CI: $11.2\%-13.4\%$), similar to the usual estimate of lifetime incidence of breast cancer of $1/8$.
(a) Posterior distribution of $\psi$.
(b) Posterior distribution of $\mu_{\textsf{prog}}$.
(c) Estimated pre-clinical hazard function.
Figure 8: Results of the analysis of the data set from the Breast Cancer Surveillance Consortium data set. Panels (a) and (b) display the marginal posterior distributions of the fraction of indolent pre-clinical cancers ($\psi$) and and the mean sojourn time among progressive prec-clinical cancers ($\mu_{\textsf{prog}}$), respectively. The posterior means (dashed red line) and $95$% credible intervals (dotted red lines) are included. Panel (c) shows the estimated pre-clinical hazard rate as a function of age, with the posterior median (solid line) and $95$% credible band (dashed lines) indicated.
6.2 Estimation of the overdiagnosis rate
Using our MCMC draws, we can estimate the posterior of any function of the parameters. Here, we focus on the extent of overdiagnosis in a screening program, defined as the the mammographic detection of cancer that would not become clinically evident in the woman’s remaining lifetime. There are two contributions to overdiagnosis. First, by definition, every non-progressive cancer that is detected on mammography is overdiagnosed. Second, progressive screen-detected cancer are overdiagnosed if the time from screen-detection to death from a breast cancer-unrelated cause is shorter than the time from screen-detection to clinical progression (also referred to as the lead-time). We use the breast cancer natural history estimates from the BCSC cohort to estimate the overdiagnosis rate in a biennial screening program from age $50$ to $74$ years. Table 3 provides a side-by-side comparison of the three models with $\alpha_{\textsf{H}}=2.5$. The three models yield similar estimate for the overdiagnosis rate and the respective contribution indolence and mortality from other sources, showing that these results are robust to the choice of $\alpha_{\textsf{prog}}$. This is reassuring given that the ALOOCV procedure was unable to delineate the best value of this parameter. Moreover, the estimates we obtain are in line with those of Ryser et al. (2022), which are based on a Markov model with a piece-wise constant onset rate for pre-clinical cancer, corroborating the claim that approximately 1 in 7 screen-detected breast cancers are overdiagnosed.
[…]
Indolence Mortality from other causes
$2.5$ $1$ $15.6$ $(5.4,28.9)$ $5.3$ $(0.1,19.2)$ $10.3$ $(2.8,22)$ $7.1$ $(5.3,9.1)$ $4.2$ $(0.1,14.9)$
$2.5$ $1.5$ $13$ $(3.8,25.7)$ $5.7$ $(0.1,19)$ $7.3$ $(1.6,17.5)$ $5.9$ $(4.5,7.6)$ $4.7$ $(0.1,15.5)$
$2.5$ $2$ $13.4$ $(3.6,26.9)$ $6.9$ $(0.2,22.3)$ $6.6$ $(1.4,15.9)$ $5.6$ $(4.1,7.1)$ $5.7$ $(0.2,17)$
Ryser et al. (2022) $15.4$ $(9.4,26.5)$ $6.1$ $(0.2,20.1)$ $9.3$ $(5.5,13.5)$ $6.6$ $(4.9,8.6)$ $4.5$ $(0.1,14.8)$
Table 3: Estimates of the overdiagnosis rate and the respective contribution of indolence and mortality from other causes for the three models with $\alpha_{\textsf{H}}=2.5$ and the model of Ryser et al. (2022) The posterior means and $95\%$ credible intervals are reported.
7 Discussion
In this paper, we introduce an efficient data-augmented MCMC sampling algorithm for fitting a flexible family of semi-Markov mixture models of cancer natural history to individual-level screening and cancer diagnosis histories. By allowing for age-dependent pre-clinical onset hazards, non-exponentially distributed sojourn times, and a mixture of indolent and progressive tumors, we substantially expand the range of cancer progression models and dynamics amenable to rigorous Bayesian inference via MCMC. These advances are a marked departure from prior approaches, which relied on simple Markov models for tractability. Our sampler efficiently explores the high-dimensional joint space of model parameters and latent variables, and automatically quantifies individual-level uncertainty of clinically relevant quantities such as a person’s pre-clinical onset time, indolence status, and sojourn time.
A key application of our methodology is the estimation of screening-related overdiagnosis rates. In the motivating example of breast cancer, overdiagnosis has been documented as the chief harm of screening, yet there is substantial residual uncertainty about its true extent. Due to the counterfactual nature of overdiagnosis, its estimation is methodologically challenging and, as in our case of observational data from a screened cohort, relies on sound estimates of the latent tumor progression dynamics. Applying our method to a high-quality cohort of women undergoing mammography screening, we make two important observations. First, our breast cancer overdiagnosis estimates are similar to those previously derived from the same cohort using a slightly different model structure, thus adding weight to the finding that approximately 1 in 7 screen-detected breast cancers are overdiagnosed. Second, although the data contain insufficient information to delineate the best-fitting sojourn time parameterization, we find that the overdiagnosis estimates were robust to changes in this parameterization. This finding of robustness is particularly important because model misspecification is guaranteed when modeling complex biological processes such as cancer progression. Finally, we note that thanks to the latent data augmentation, our approach provides individual-level estimates of the time of onset with pre-clinical cancer. These data could, for example, be leveraged to derive individual level estimation of overdiagnosis.
Future efforts will be directed toward methodological extensions of our approach, including covariate-dependent hazard rates, non-parametric formulations of the pre-clinical onset hazard for additional flexibility, and the introduction of early and late disease stages. In the meantime, the current iteration of the method, as available through the R package baclava, has the potential to facilitate rigorous natural history estimation across various cancers with available screening data.
Acknowledgements.
This research was funded by the National Cancer Institute (R35-CA274442) and and the National Science Foundation (DMS 2230074 and PIPP 2200047). Data collection for this research was funded by the National Cancer Institute (P01CA154292, U01CA63736, U01CA069976). The collection of cancer and vital status data was supported in part by several state public health departments and cancer registries (https://www.bcsc-research.org/work/acknowledgement). We thank the participating women, mammography facilities, and radiologists for the data they have provided. You can learn more about the BCSC at: http://www.bcsc-research.org/. All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the National Cancer Institute or the National Institutes of Health.
Appendix A Proof of Theorem 1
Proof.
From Equations (9), (13) and (18), the joint distribution of $(\theta,\psi,\beta,\mathbf{Z},\mathbf{Y})$ is
$\pi(\theta,\psi,\beta,\mathbf{Z},\mathbf{Y})$ $=\pi(\theta,\psi,\beta)f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})L(\theta,\psi;\mathbf{Z})$
$=\pi(\beta)\pi(\theta,\psi)\beta^{m^{+}}(1-\beta)^{m^{-}}\prod_{i=1}^{n}f_{t}\left(t^{PC}_{i}\mid Z_{i},\theta\right)L(\theta,\psi;\mathbf{Z}),$ (47)
From Equation (A), the full conditional distribution of $\beta$ is, by proportionality,
$\pi(\beta\mid\theta,\psi,\mathbf{Z},\mathbf{Y})$ $\propto\pi(\theta,\psi,\beta,\mathbf{Z},\mathbf{Y})$
$\propto\pi(\beta)\beta^{m^{+}}(1-\beta)^{m^{-}}$
$\propto\beta^{a+m^{+}-1}(1-\beta)^{b+m^{-}-1},$
which is the kernel of the distribution $\textsf{Be}\left(a+m^{+},b+m^{-}\right)$.
Appendix B Metropolis-Hastings updates for the Weibull rate parameters of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$
The proposal distributions in the Metropolis-Hastings steps for $\lambda_{\textsf{H}}$ and $\lambda_{\textsf{prog}}$ correspond to normal random walks on the positive line with a reflection at 00. At step $m=1,\dots,M$, the candidate values are
$\lambda_{\textsf{H}}^{\star}\mid\lambda_{\textsf{H}}^{(m-1)}=|\lambda_{\textsf{H}}^{(m-1)}+\eta_{\textsf{H}}|,\quad\eta_{\textsf{H}}\sim N(0,\epsilon_{\textsf{H}}),$ (48)
and
$\lambda_{\textsf{prog}}^{\star}\mid\lambda_{\textsf{prog}}^{(m-1)}=|\lambda_{\textsf{prog}}^{(m-1)}+\eta_{\textsf{prog}}|,\quad\eta_{\textsf{prog}}\sim N(0,\epsilon_{\textsf{prog}}),$ (49)
where $\epsilon_{H}$ and $\epsilon_{P}$ are tuning parameters and the absolute values ensure that the proposed values are positive.
Simple algebra shows that the proposal distributions (48) and (49) are symmetrical. For $\lambda_{\textsf{H}}$, the Metropolis-Hastings acceptance probability (20), therefore, simplifies to
$\alpha\left(\lambda_{\textsf{H}}^{m-1}\rightarrow\lambda_{\textsf{H}}^{\star}\right)=\min\left\{1,\dfrac{\pi(\theta^{\star},\psi,\beta,\mathbf{Z}\mid\mathbf{Y})}{\pi(\theta^{(m-1)},\psi,\beta,\mathbf{Z}\mid\mathbf{Y})}\right\}=\min\left\{1,\dfrac{\pi(\theta^{\star})L(\theta^{\star},\psi;\mathbf{Z})}{\pi(\theta^{\star})L(\theta^{\star},\psi;\mathbf{Z})}\right\},$
where $\theta^{\star}=\left(\lambda_{\textsf{H}}^{\star},\lambda_{\textsf{prog}},\psi\right)$ contains the proposed value for $\lambda_{\textsf{H}}$, and $\theta^{(m-1)}=\left(\lambda_{\textsf{H}}^{(m-1)},\lambda_{\textsf{prog}},\psi\right)$ the current value. The second equality holds by Equations (9)-(17). The acceptance probability for $\lambda_{\textsf{prog}}$ is analogous. We use adaptive Markov chain Monte Carlo to find values of $\epsilon_{\textsf{H}}$ and $\epsilon_{\textsf{prog}}$ that yield a pre-specified acceptance rate; details are provided in Appendix E.
Appendix C Hypothesis test for approximate leave-one out cross-validation
Given two models $A$ and $B$, the model with the largest predictive fit (36) is better. Consider the difference between the estimates of the two models
$\Delta_{AB}=\widehat{\textsf{pf}}_{A}-\widehat{\textsf{pf}}_{B}=\sum_{i=1}^{n}\Delta_{AB}^{i},$
where $\Delta_{AB}^{i}=\log(\hat{f_{A}}(y_{i}\mid\mathbf{Y}_{-i}))-\log(\hat{f_{B}}(y_{i}\mid\mathbf{Y}_{-i}))$ is the difference between the contributions of individual $i$. We now describe a paired t-test for the null hypothesis that $\Delta_{AB}=0$. To estimate the the standard error of $\Delta_{AB}$, we take advantage of the fact that the data of the same subjects are used to compute $\hat{pf}_{A}$ and $\hat{pf}_{B}$ and use the paired estimate
$SE(\Delta_{AB})=\sqrt{n\textsf{Var}_{i=1}^{n}(\Delta_{AB}^{i})},$ (50)
where $\textsf{Var}_{i=1}^{n}a_{i}=\frac{1}{n-1}\sum_{i=1}^{n}(a_{i}-\bar{a})^{2}$.
We can then appeal to the Central Limit Theorem to describe the asymptotic behavior of $\Delta_{AB}$. By exchangeability of the subjects, $\Delta_{AB}$ is a sum of identically distributed random variables with finite variance. These random variable are, however, not entirely independent; the source of dependence between $\Delta_{AB}^{i}$ and $\Delta_{AB}^{j}$ comes from conditioning on the data of subject $j$ when evaluating $\Delta_{AB}^{i}$ and conditioning on the data of subject $i$ from $j$ when evaluating $\Delta_{AB}^{j}$. For large $n$, however, the influence of a single subject becomes negligible, resulting in quasi independence between $\Delta_{AB}^{i}$ and $\Delta_{AB}^{j}$. The conditions for a paired $t$-test are, therefore, satisfied in large samples, with the $t$-statistic
$t=\frac{\Delta_{AB}}{SE(\Delta_{AB})},$ (51)
and $n-1$ degrees of freedom.
Appendix D Details of the simulated data in Section 5
The age of the first screen is sampled from the categorical distribution with weights $w_{k}\propto\exp\{-k/5\}$ for $k=40,41,\dots,80$. Inter-screen intervals are sampled from the distribution $\textsf{Poisson}(0.5)$ plus $1$ year, resulting in a mean inter-screen interval of $1.5$ years. Finally, for individual $i=1,\dots,n$, the age of last follow-up is $\min\{100,t^{i}_{1}+W_{i}\}$ with $W_{i}\sim\textsf{Exp}(1/5)$ independently over $i$, implying that individuals remain in the screening program for $5$ years on average.
Appendix E Adaptive Markov chain Monte Carlo
The Metropolis-Hastings steps for $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$, $\lambda_{\textsf{H}}$ and $\lambda_{\textsf{prog}}$ rely on normal random walk proposals governed by a step size parameter. We use the adaptive procedure of Hoffman et al. [2014] for finding values for the step size parameters that result in a pre-specified acceptance rate. We now describe this procedure.
Consider the generic case of a random walk Metropolis-Hastings step where the normally distributed increments have variance $\epsilon>0$. $\epsilon$ acts as a step size parameter, and we wish to find a value $\bar{\epsilon}$ such that the acceptance rate reaches some desired value $\delta\in(0,1)$.
Write $H_{m}=\delta-\alpha_{m}$, where $\alpha_{m}\in[0,1]$ is the Metropolis-Hastings acceptance probability at step $m=1,\dots,M$. We apply the update
$\log(\epsilon_{m+1})=(1-\omega_{t})\log(\epsilon_{t})+\omega_{m}\xi_{m+1},$ (52)
where
$\xi_{m+1}=-\frac{\sqrt{m}}{\gamma}\frac{1}{m+m_{0}}\sum_{i=1}^{m}H_{m},$ (53)
is the correction term at step $m+1$ and $\omega_{m}=m^{-\kappa}$.
This iterative procedure is a stochastic approximation method. Formula (52), introduced by Hoffman et al. [2014], gives more weight to later iterations than the classical update of Robbins and Monro [1951]. This is desirable in the setting of MCMC as early iterations correspond to the transient phase of the Markov chain which may be quite different from its recurrent phase. Formula (52), therefore, ensures that the early iterations do not have a disproportionate weight in the tuning procedure.
In the Metropolis-Hastings step for $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$, we set the desired acceptance rate to $0.24$ as it is a high-dimensional update, while in the steps for $\lambda_{\textsf{H}}$ and $\lambda_{\textsf{prog}}$ we set it to $0.44$ as they are univariate updates [Gelman et al., 1996]. Following Hoffman et al. [2014], we further set $\gamma=0.05$, $m_{0}=10$, and $\kappa=0.75$, which is sufficient to ensure that $\sum_{m=1}^{\infty}\omega_{m}=\infty$ and $\sum_{m=1}^{\infty}\omega_{m}^{2}<\infty$, as is required for $\epsilon_{m}$ to converge to $\bar{\epsilon}$ [Andrieu and Thoms, 2008].
Given a total number of iteration $M$ and a number of warm up iterations $W<M$, we use the iterations $m=1,\dots,W$ as a warm up phase during which the step size $\epsilon$ is tuned using Formula (52). We then use the value $\epsilon_{W}$ found at the end of the warm-up phase for the remaining iterations. The warm-up iterations are discarded when we estimate the posterior distribution.